OpenStudy (anonymous):

i need help please? Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (anonymous):

i guess it is pretty simple!

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) y + 1= 6 ( x-5)

OpenStudy (anonymous):

when a number is negative since subtraction is in the equation you add instead

OpenStudy (anonymous):

you are given that m=(slope)=6 right? y1=-1 and x1=5 plugging in the value in the equation y – y1 = m(x – x1) you get y+1=6(x-5) ie y+1=6x-30 y+31-6x=0 HENCE THE ANSWER FOR THE EQUATION FOR THE LINE IS y-6x+31=0

OpenStudy (anonymous):

hope i helped! #rawr im a tiger cheers!

OpenStudy (anonymous):

rawr shes angry run away. hahahaha.